OpenStudy (bananas1234):

Simplify: (3y)^4 times y^-2

Nnesha (nnesha):

\[\huge\rm x^{-m} = \frac{ 1 }{ x^m }\]

OpenStudy (bananas1234):

i'm still kinda confused, could you walk me through the steps?

Nnesha (nnesha):

sure i was afk okay so whenever you have NEGATIVE exponent you should change that to its reciprocal for example \[\huge\rm x^{-m} = \frac{ 1 }{ x^m }\]

OpenStudy (bananas1234):

ohhh i see so y^-2 would be 1/y-2?

OpenStudy (bananas1234):

or 1/y^2?

Nnesha (nnesha):

nope ......look at the exponent rules why do you have to flip that bec to change the negative sign to positive

Nnesha (nnesha):

yes that's right

OpenStudy (bananas1234):

ok

Nnesha (nnesha):

now (3y)^4 solve this \[\huge\rm (3y)^4 \times \frac{ 1 }{ y^2 }\] what is (3y)^4= ?

OpenStudy (bananas1234):

um 81y times 1/y^2

Nnesha (nnesha):

81 and what ??

OpenStudy (bananas1234):

im not sure maybe 81y^2?

Nnesha (nnesha):

not sure ?? >.^ look at the previous post we just did one question same like this

OpenStudy (bananas1234):

1/81y^2?

Nnesha (nnesha):

nope just the parentheses part (3y)^4 = ?

OpenStudy (bananas1234):

81y?

Nnesha (nnesha):

how did you get 81 ? explain

OpenStudy (bananas1234):



Nnesha (nnesha):

yes bec there is 4th power of 3 what about y y is also in parentheses .....

OpenStudy (bananas1234):

i dont know what to do with the y

Nnesha (nnesha):

how did we solve last question exponent rule (x^m)^n = x^{ m times n}

Nnesha (nnesha):

y there is an exponent of y which is one \[(y^1)^4\]

OpenStudy (bananas1234):

oh so 81y^1? or would i add the exponents getting 3y^5

Nnesha (nnesha):

 \[\huge\rm (x^m)^n = x^{m \times n }\] multiply inside exponent of variable by outside one